OpenStudy (anonymous):

A certain box of cereal should have a net weight that is with 1.5 grams of 510. After the first phase of production the weight of the cereal is 308 grams. By how much should the weight of the cereal increase to remain within the specified weight? A. Between 200.5 and 203.5 grams B. between 201.5 and 204.5 grams C. Between 202.5 and 203.5 grams D. Between 201.5 and 203.5 grams

OpenStudy (shinalcantara):

Since the range for the net weight of the cereal is 510(+-) 1.5 grams, then you can have the net weight as 510 + 1.5 = 511.5 grams to 510 - 1.5 = 508.5 grams ---------------------------- Since in the first production, the weight is 308 grams, then the weight that should be added to comply with the limits would be 511.5 - 308 = 203.5 grams 508.5 - 308 = 200.5 grams --------------------------- Therefore it should be increased by 200.5 to 203.5 grams